OpenStudy (anonymous):

I need help with Algebra II equation. This is the third time I'm asking and I still didn't get help. It's confusing so please help.

OpenStudy (anonymous):

@grray7

OpenStudy (anonymous):

@radar

Nnesha (nnesha):

ok i'm tagging someone for you @terenzreignz first time i saw him in this website i don't know how he gonna react :(

OpenStudy (anonymous):

Thanks x

terenzreignz (terenzreignz):

hmmn...

terenzreignz (terenzreignz):

Well, a question; how were you taught how to that?

OpenStudy (anonymous):

How to do the problem that I don't know how to solve? Well, k12 slides. It doesn't make sense so then I decided to surf the internet on how to solve that... it still was confusing.

OpenStudy (anonymous):

What I would do is: For any one of the equations, solve for a variable: it could be x = (something with y and z) or, y = (something with x and z) or z, = (something with x and y) Take that and plug this every time you see the variable in the other 2 equations You should have a system of 2 variables with 2 equations. Then solve from there.

OpenStudy (compassionate):

The strategy is to reduce this to two equations in two unknowns. Do that by eliminating one of the unknowns from two pairs of equations: either from equations 1) and 2), or 1) and 3), or 2) and 3). \[3x - 2y + 2z = 30 \]\[-x+3y - 4z = -33\]\[2x-4y+3z=42\] You can start by elimating x, for example.

OpenStudy (anonymous):

-3?

OpenStudy (compassionate):

What did you eliminate from? 1) and 2), or 1) and 3), or 2) and 3)?

OpenStudy (anonymous):

Uh.... 1 and 2

OpenStudy (compassionate):

Your goal is to eliminate x. You eliminated everything. Take 1 and 2, and solve it like a normal two-system equation. Eliminate either x, y, or z. It doesn't matter.

Nnesha (nnesha):

quick question do we have to do all of 3 eliminate together or we can select two equation first ??

OpenStudy (anonymous):

I took the first and the second equation . What do i do next?

OpenStudy (compassionate):

Now Solve the system for x, y or z. Do you know how?

OpenStudy (anonymous):

I don't know how to solve this equation at all. I suck at math and worst of all... I'm in an online school where I don't have a physical teacher that I can ask for help and get a response right away.

OpenStudy (compassionate):

Alright. Well, let me explain to you what a System of Equation is. Basically, a System of Equations are like Linear Equations, except you're solving two equations. An example of a linear equation is: ilovchuu paid Compassionate 9 gold fish total. For tummy tickles, Compassionate gets 1 goldfish, and for Satanic Blood Rituals, Compassionate gets 3 whole gold fish. How many gold fish does Compassionate have from tummy tickles and Satanic Blood Rituals? The Linear equation would be: 1x + 3y = 9 Then you can either solve for x or y to find either value. Yes? Are you following so far?

OpenStudy (anonymous):

yes I am

OpenStudy (compassionate):

Whoops. Got distracted. Hehe. Sorry! Okay, so sometimes you will come across problems that utilize TWO linear equations. This is what is known as a System of Equations (System = more than 1). There are two ways to solve a system. Either by substitution or elimination. Lets say I'm asked to solve 2x + = 4 x = 4 Okay, here, I am always given x. I can simply use the "Substitution method" to solve. I simply substitute x into the first equation, and walla, I solve. Now, elimination is different. This is kinda like canceling. For example, if I'm given: 2x + 3y = 5 -1x + 3y = 4 I need to say, "what variable do I want to solve by? What would be easiest? Hmm... I can solve for either x, y or z... Lets solve for x. Now, what do I need to multiple either one of my equations by to get them to cancel? Well, if I multiply by bottom equation by 2, I can cancel my x's because 2x and -2x will cancel out." So if I have 2x + 3y = 4 -2x + 6y = 8 (Upon multiplying EVERYTHING by 2) You'll see the two's will cancel, leaving you with Then you will simply add the rest of the terms together. So 3y + 6y and 4 + 8 9y = 12 Simply divide and you have your y value. Then take your y value, plug it into EITHER equation, then you get your x value. Bim bam, done! Take a look at this example: https://www.khanacademy.org/math/algebra/systems-of-eq-and-ineq/fast-systems-of-equations/v/solving-systems-of-equations-by-elimination [Watch all of it so you can understand. Solving 3 systems is basically the same with an extra step]

OpenStudy (compassionate):

Once you understand simple two-systems, we can move onto 3-systems.

OpenStudy (midhun.madhu1987):